OpenStudy (anonymous):

what is the integral of x/(1+x) dx ?

OpenStudy (amistre64):

If we add 1 and subtract 1 up top, what does that simplify to?

OpenStudy (amistre64):

\[\int \frac{x}{x+1}dx\] \[\int \frac{x+1-1}{x+1}dx\] \[\int \frac{x+1}{x+1}-\frac{1}{x+1}dx\] \[\int 1-\frac{1}{x+1}dx\]